OpenStudy (anonymous):

Find the domain of the exponential function f(x) = -2(3)^x. A. Any real numbers B. Positive real numbers C.Negative real numbers D. Non-zero real numbers

OpenStudy (tkhunny):

Domain... Is that the values that go in or the values that come out?

OpenStudy (anonymous):

So would it be b then?

OpenStudy (tkhunny):

Why is there a restriction at all?

OpenStudy (anonymous):

Because i dont know how to do these types of questions

OpenStudy (tkhunny):

Positive exponents are okay. Negative exponents are okay. Zero exponents are okay.

OpenStudy (tkhunny):

Generally, it is an elimination process for the Domain. As a general rule, just start with "All Real Numbers" and see if you need to eliminate anything,

OpenStudy (tkhunny):

A few quick examples: f(x) = e^x -- No restriction g(x) = 1/(x-2) -- Well, we'd better not try x = 2 h(x) = log(x) -- By definition, x > 0 r(x) = 2x+5 -- No restriction. s(x) = sqrt(x) -- x >= 0